A rescue plane is flying horizontally at a height of 97 m above ground. The pilot spots a survivor and releases an emergency kit with a parachute. The kit descends at a constant vertical acceleration of -6.63 m/s2 and the initial plane horizontal speed is 66.4 m/s. Assuming air resistance and wind are negligible, how far away from the survivor will the emergency kit hit the ground?


1
Expert's answer
2021-04-19T17:11:13-0400

Find the time it will take to reach the ground at -6.63 m/s/s:


"t=\sqrt{\frac{2h}{a}}."

Find the distance the horizontal kit will cover during this time:


"r=vt=v\sqrt{\frac{2h}{a}}=359\text{ m}."